### v1

# Some examples of use of transfinite induction in analysis

Abstract
It is not uncommon in analysis that existence of extremal objects is obtained via an iterative procedure: we start from a given admissible object, then modify it, then modify again etc… If being extremal means maximimizing a real valued quantity and we are sure to approach the supremum fast enough, after a countable number of steps and a limiting procedure we are done.
In this short note we want to advertise a slightly different line of thought, where rather than trying to approach the supremum fast enough, we: try to increase, if possible, the function to be maximized and, at the same time, index our recursive procedure over ordinals. Since there are no increasing functions from $\omega_{1}$ to $\mathbb{R}$, the procedure must stop at some countable ordinal and existence is proved anyway.
The advantage of this line of reasoning is that it can be helpful even in situations where it is not so evident how to measure ‘being maximal’ via a real valued function. This is the case, for instance, for existence of a Maximal Globally Hyperbolic Development of an initial data set in General Relativity.
Speaking of this particular example, we also show that such ‘real-valued quantification’ of the size of a development is actually possible, thus existence of a maximal one can be obtained in a countable number of steps using the original argument in [2] together with the standard procedure depicted above. This provides a way alternative to the one given in [5] to ‘dezornify’ the proof in [2].
1 Introduction
In many situations in analysis we want to prove existence of a certain extremal object, but perhaps such existence is not immediately evident, and we must proceed by approximation. If being extremal means maximizing a real valued function, then a typical procedure would be: start from a given admissible object, then modify it by trying to approach the supremum, then modify it again etc… If we are sure to approach the supremum fast enough, after a countable number of steps and a limiting procedure we are done. We shall informally call this ‘proof via big steps’.
In this paper we want to advertise a different kind of argument, that somehow replaces these ‘big steps’ with possibly ‘small ones’. It relies on the fact that any monotone map from the first uncountable ordinal to the reals must be eventually constant, thus if we index our iterative procedure over the ordinals and at every step we try to augment our given real valued function, then regardless of how big our steps are, we are sure that after a countable number of steps we are done anyway.
We shall illustrate this principle with three practical applications, presented in a somehow increasing order of difficulty.
There is a clear downside in our approach, namely the fact that it relies on concepts such as ordinal numbers and transfinite recursion, that are rarely found in analysis’ papers. The upside, though, is that it might work even in situations where it is not really clear ‘how to quantify progresses’ in our construction. This is the case, in particular , for existence of a the Maximal Global Hyperbolic Development of an initial data set, and in this sense it is no surprise that the original proof in [2] relied on the Axiom of Choice. In this setting, the key observation that allows to reduce the number of Choices to a countable quantity is the statement by Geroch [3]:
A smooth connected manifold carrying a non-degenerate metric tensor is separable (1.1)
whose proof we shall report in Proposition 3.4. We think at this statement as: a non-degenerate metric tensor gives, regardless its signature, a ‘scale measured by reals’ to the given manifold, and once such a scale is present, the object must be separable.
With all this being said, we shall also provide an alternative proof of the existence of a Maximal Globally Hyperbolic Development that has nothing to do with transfinite recursion and ordinals. Rather, we construct a way of quantifying the size of a development, so that a maximal one can be built via ‘big steps’, along classical lines. Still, the fact that this quantification is not-so-evidently found is perhaps an argument in favour of the relevance of the line of reasoning discussed here in the mathematical conversation.
I thought for the first time at a ‘small steps’ argument via transfinite recursion of the kind just discussed in my PhD thesis [4, Proposition 5.10]. I mentioned this to a number of colleagues and to all of them the argument seemed new; still, I cannot exclude that in circles far from mine this is a well-known line of reasoning. In [4] I needed a statement closely resembling Ekeland’s variational principle, of which I was unaware, and admittedly I did not find a way to make a proper quantification, thus I resorted on this other approach. Later, any time I was tempted to use this line of thought, a quantification was easily found, making the small-steps argument irrelevant. I thus never published anything on the matter, limiting myself to asking questions on the matter on MathOverflow [1], especially in connection to the kind of Axiom of Choice needed to carry out the argument.
I realized only recently the relevance of this line of thought in relation to Maximal Global Hyperbolic Developments, and stimulated by the above considerations looked for a way to measure the size of developments via a real valued function.
Acknowledgment I wish to warmly thank Prof. A. Karagila for a series of conversations that helped me understanding the equivalence of ${\rm DC}_{\omega_{1}}$ and Lemma 2.1.
This study was funded by the European Union - NextGenerationEU, in the framework of the PRIN Project Contemporary perspectives on geometry and gravity (code 2022JJ8KER – CUP G53D23001810006). The views and opinions expressed are solely those of the author and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
2 Abstract presentation of the mechanism
The next lemma gives an abstract presentation of the argument we are going to use:
Lemma 2.1 (Transfinite recursion over countable ordinals).
Let $A$ be a set, $F\subset A$ a subset and $\mathcal{S}$ a non-empty collection of sequences of length $\omega_{1}$ with the following properties:
  • i)
    (if $F$ has not been reached, then we can continue)
    If $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$ is so that $a_{\alpha^{\prime}}\notin F$ for every $\alpha^{\prime}<\alpha$, then there is $a_{\alpha}\in A$ such that $(a_{\alpha^{\prime}})_{\alpha^{\prime}\leq\alpha}\in\mathcal{S}$;
  • ii)
    (limits of admissible sequences are admissible)
    If $\alpha<\omega_{1}$ is a limit ordinal and $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}$ is so that $(a_{\alpha^{\prime\prime}})_{\alpha^{\prime\prime}<\alpha^{\prime}}\in\mathcal{S}$ for every $\alpha^{\prime}<\alpha$, then $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$;
  • iii)
    (there is no admissible sequence $\omega_{1}$-long)
    There is no $(a_{\alpha})_{\alpha<\omega_{1}}$ such that for every $\alpha<\omega$ we have $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$.
Then there is $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$ and $\bar{\alpha}<\alpha$ so that $a_{\bar{\alpha}}\in F$. In particular, $F$ is not empty.
Proof.
Define recursively a transfinite sequence as follows. Let $a_{0}\in A$ be arbitrary. Then let $\alpha<\omega_{1}$ and assume to have already defined $a_{\alpha^{\prime}}$ for every $\alpha^{\prime}<\alpha$} and that $(a_{\alpha^{\prime\prime}})_{\alpha^{\prime\prime}\leq\alpha^{\prime}}\in\mathcal{S}$ for every $\alpha^{\prime}<\alpha$. Then $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$: this is obvious if $\alpha$ is a successor, while if it is limit we use $(ii)$. If for some $\alpha^{\prime}<\alpha$ we have $a_{\alpha^{\prime}}\in F$ we are done, otherwise by $(i)$ there is $a_{\alpha}\in A$ such that $(a_{\alpha^{\prime}})_{\alpha^{\prime}\leq\alpha}\in\mathcal{S}$.
Proceed by transfinite recursion until either we find an $\omega_{1}$-long sequence $(a_{\alpha})_{\alpha<\omega_{1}}$ such that for every $\alpha<\omega_{1}$ we have $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$ or we find some $\alpha<\omega_{1}$ and $\bar{\alpha}<\alpha$ such that $(a_{\alpha^{\prime}})_{\alpha^{\prime}<\alpha}\in\mathcal{S}$ so that $a_{\bar{\alpha}}\in F$. Assumption $(iii)$ tells that the former case does not happen, thus the latter holds. ∎
We might replace assumption $(iii)$ above with the assumption that sequences in $\mathcal{S}$ are injective, as this is what happens in the applications we have in mind. However, for us it might be relevant to know that the recursive procedure only involves countable ordinals (e.g. in discussing the application to the Hahn-Jordan decomposition we need to be sure that the sets we produce at each step are measurable) and for this assumption $(iii)$ matters. A further advantage of having it is that it clarifies which version of Axiom of Choice is needed to carry on the argument: as Prof. A. Karagila patiently explained me, Lemma 2.1 is equivalent to ${\rm DC}_{\omega_{1}}$.
3 Three examples
3.1 The Hahn-Jordan decomposition
The Hahn-Jordan decomposition is a basic statement in measure theory telling that a finite signed measure $\mu$ can uniquely be written as difference of two non-negative signed measures concentrated on disjoint subsets. Its proof relies on the following lemma. Recall that given a finite signed measure $\mu$, a measurable set is called ‘negative’ if it has no measurable subset with positive measure.
Lemma 3.1.
Let $\mu$ be a finite signed measure on the measurable space $({\rm X},\mathcal{A})$ and $E\in\mathcal{A}$.
Then there is $E^{\prime}\subset E$, $E\in\mathcal{A}$, negative with $\mu(E^{\prime})\leq\mu(E)$.
Proof via ‘big steps’ For every integer $n\in\mathbb{N}$ we define a set $E_{n}\in\mathcal{A}$ as follows. Let $E_{0}:=E$ and recursively define $t_{n}:=\sup_{C\subset E_{n}}\mu(C)$, then pick $C_{n}\subset E_{n}$ so that $\mu(C_{n})\geq\min\{1,\tfrac{t_{n}}{2}\}$ and let $E_{n+1}:=E_{n}\setminus C_{n}$.
Put $E_{\infty}:=\cap_{n}E_{n}=E\setminus\cup_{n}C_{n}$ and notice that $\mu(E_{\infty})=\mu(E)-\sum_{n}\mu(C_{n})\leq\mu(E)$. Suppose $E_{\infty}$ is not negative. Then there is $C\subset E_{\infty}\subset E_{n}$ for every $n$ with $\mu(C)>0$. Thus $t_{n}>\mu(C)$ for every $n$ and therefore $\mu(C_{n})\geq\min\{1,\tfrac{\mu(C)}{2}\}$. Since the $C_{n}$’s are disjoint we would get $\mu(\cup_{n}C_{n})=\sum_{n}\mu(C_{n})=+\infty$, contradicting the fact that $\mu$ is finite. $\square$
Proof via ‘small steps’ We shall use Lemma 2.1 with $A:=\mathcal{A}$, $F\subset A$ the collection of negative subsets $E^{\prime}$ of $E$ with $\mu(E^{\prime})\leq\mu(E)$ (equivalently: of those subsets $E^{\prime}$ of $E$ so that for no $E^{\prime\prime}\subset E^{\prime}$ we have $\mu(E^{\prime\prime})<\mu(E^{\prime})$) and $\mathcal{S}$ the collection of decreasing sequences $(E_{\beta})_{\beta<\alpha}$ of subsets of $E$ with $E_{0}:=E$ and $\beta\mapsto\mu(E_{\beta})$ strictly decreasing. The fact that these satisfy properties $(i),(ii)$ in Lemma 2.1 obvious. Property $(iii)$ follows from the fact that there are no strictly increasing maps from $\omega_{1}$ to $\mathbb{R}$.
Lemma 2.1 grants that $F$ is not empty, which is the claim. $\square$
3.2 Ekeland’s variational principle
The variational principle of Ekeland allows to show existence of almost minimizers of a lower semicontinuous functional in absence of compactness (but in presence of completeness).
A possible formulation is the following:
Theorem 3.2.
Let $({\rm X},{\sf d})$ be a complete metric space, $f:{\rm X}\to[0,+\infty]$ be lower semicontinuous and $\bar{x}\in{\rm X}$. Then there is $\bar{y}\in{\rm X}$ with $f(\bar{y})+{\sf d}(\bar{x},\bar{y})\leq f(\bar{x})$ such that
$f(z)+{\sf d}(z,\bar{y})\geq f(\bar{y})\qquad\forall z\in{\rm X}.$ (3.1)
Proof via ‘big steps’ Define the relation $\leq$ on $\rm X$} by declaring that $z_{1}\leq z_{2}$ whenever
$f(z_{1})+{\sf d}(z_{1},z_{2})\leq f(z_{2}).$
It is clear that this is a partial order and by lower semicontinuity of $f$ that $\{y\in{\rm X}:y\leq x\}$ is closed for every $x\in{\rm X}$. Now consider the given $\bar{x}$ and notice that if there is no $x\leq\bar{x}$ with $f(x)<\infty$, then the choice $\bar{y}:=\bar{x}$ satisfies the conclusion, otherwise replacing $\bar{x}$ with such $x$ we can assume that $f(\bar{x})<\infty$.
For every $x\in{\rm X}$ define
$i(x):=\inf_{y\leq x}f(y),\qquad\text{ and }\qquad D(x):=f(x)-i(x)\geq 0$
and recursively define a sequence $(x_{n})\subset{\rm X}$ by putting $x_{0}:=\bar{x}$ and then, given $x_{n}$, find $x_{n+1}\leq x_{n}$ with $f(x_{n+1})\leq\tfrac{1}{2}(i(x_{n})+f(x_{n}))$. It is clear that such $x_{n+1}$ can be found, that ${\sf d}(x_{n+1},x_{n})\leq D(x_{n})$ and that $D(x_{n+1})\leq\tfrac{1}{2}D(x_{n})$. Since $D(x_{0})<\infty$ we conclude that $(x_{n})$ is Cauchy, hence it converges to a limit $x_{\infty}$. The lower semicontinuity of $f$ and the fact that $i$ is $\leq$-non-decreasing show that $D(x_{\infty})=0$. This means that $f(x_{\infty})=i(x_{\infty})$ and thus that $\bar{y}:=x_{\infty}\leq\bar{x}$ is the desired point. $\square$
Proof via ‘small steps’ Define the relation $\leq$ on $\rm X$} by declaring that $z_{1}\leq z_{2}$ whenever
$f(z_{1})+{\sf d}(z_{1},z_{2})\leq f(z_{2}).$
It is clear that this is a partial order and by lower semicontinuity of $f$ that $\{y\in{\rm X}:y\leq x\}$ is closed for every $x\in{\rm X}$. Now consider the given $\bar{x}$ and notice that if there is no $x\leq\bar{x}$ with $f(x)<\infty$, then the choice $\bar{y}:=\bar{x}$ satisfies the conclusion, otherwise replacing $\bar{x}$ with such $x$ we can assume that $f(\bar{x})<\infty$.
We use Lemma 2.1 with $A:={\rm X}$, $F\subset A$ the collection of points $\leq\bar{x}$ that are $\leq$-minimizers, i.e. points $x$ so that $y\leq x$ implies $y=x$, and $\mathcal{S}$ the collection of continuous, injective and $\leq$-non-increasing sequences bounded from above by $\bar{x}$. It is clear that assumption $(ii)$ of Lemma 2.1 are satisfied. $(iii)$ is satisfied as well because along any sequence in $\mathcal{S}$ the function $f$ is real valued, monotone and injective, and we already recalled that no such map can be defined on $\omega_{1}$. It thus remains to show $(i)$. This is obvious if $\alpha$ is a successor ordinal, as in this case any $\alpha$-sequence $(x_{\beta})_{\beta<\alpha}$ has $x_{\alpha-1}$ as last element, and if it is not $\leq$-minimizer, an $x_{\alpha}\lneq x_{\alpha-1}$ exists. Thus say that $\alpha$ is a limit ordinal. We claim that for every $\alpha$-sequence $(x_{\beta})_{\beta<\alpha}\in\mathcal{S}$ we have
$f(x_{\beta})+{\sf d}(x_{\beta},x_{\gamma})\leq f(x_{\gamma})\qquad\forall\gamma\leq\beta<\alpha.$ (3.2)
We show this by transfinite induction on $\beta$ ($\alpha$ is fixed here and can be assumed to be >0>0). It is clear that (3.2) holds for $\beta=0$ and that its validity for $\beta$ implies that for $\beta+1$. For $\beta$ limit, let $(\beta_{n})$ be an increasing sequence whose supremum is $\beta$ and notice that (3.2) and the inductive assumption show that $n\mapsto x_{\beta_{n}}$ is a Cauchy sequence: its limit must, by continuity of $(x_{\cdot})$, coincide with $x_{\beta}$. Also, by lower semicontinuity of $f$ we see that (3.2) holds for $\beta$.
Thus (3.2) holds. Now let $(\alpha_{n})$ be increasing with supremum $\alpha$. As before, (3.2) implies that $n\mapsto x_{\alpha_{n}}$ is Cauchy, hence admits a limit and this limit does not depend on the sequence $(\alpha_{n})$ chosen, as interlacing any two such sequences we still have that the limit exists. Call $x_{\alpha}$ such limit and notice that, by construction and arguments already used, the sequence $(x_{\beta})_{\beta\leq\alpha}$ belongs to $\mathcal{S}$, as desired.
It follows by Lemma 2.1 that $F$ is not empty, which was the claim. $\square$
3.3 Maximal Globally Hyperbolic Development
An initial data set is a triple $(\Sigma,h,\kappa)$ with $\Sigma$ being a smooth 3-dimensional manifold, $h$ a smooth Riemannian metric on it and $\kappa$ a symmetric 2-form satisfying the constraint equations
$\begin{split}R+({\rm tr}_{h}\kappa)^{2}-\|\kappa\|^{2}_{h}&=0,\\ {\rm div}_{h}(\kappa)-{\rm d}({\rm tr}_{h}\kappa)&=0.\end{split}$
A development of $(\Sigma,h,\kappa)$ is a triple $(M,g,\varphi)$ where:
  • i)
    $(M,g)$ is a Ricci-flat spacetime (and thus solves Einstein’s equations in the vacuum),
  • ii)
    $\varphi:\Sigma\to M$ is a smooth isometric diffeomorphism of $\Sigma$ and $S:=\varphi(\Sigma)\subset M$,
  • iii)
    $\varphi$ sends $\kappa$ to the second fundamental form of $S$,
  • iv)
    $S$ is a Cauchy hypersurface of $M$.
In particular, $M$ admits a Cauchy hypersurface and thus is globally hyperbolic. For the terminology we refer to [2]. We observe that the embedding $\varphi:\Sigma\to M$ induces, for every $p\in\Sigma$, a unique causal isometry $\Phi_{p}$ from $(\mathbb{R}\times T_{p}\Sigma,{\rm d}t^{2}-h)$ to $(T_{\varphi(p)}M,g)$ sending $(0,v)$ to ${\rm d}\varphi(v)$, being intended that $(1,0)\in\mathbb{R}\times T_{p}\Sigma$ is future directed.
[…]
${\rm d}\psi_{\varphi(p)}\circ\Phi_{p}=\tilde{\Phi}_{p}\qquad\forall p\in\Sigma.$ (3.3)
We discuss uniqueness. The fact that any smooth causal isometry $\psi:M\to\tilde{M}$ must satisfy (3.3) is obvious, and so is the fact that (3.3) fully characterizes ${\rm d}\psi:T_{\varphi_{p}}M\to T_{\tilde{\varphi}(p)}\tilde{M}$. Now for uniqueness of $\psi$ let $q\in M$ and $\gamma$ an inextendible timelike geodesic (intended as solution of the geodesic equation) passing through $q$. Since $\varphi(\Sigma)$ is a Cauchy hypersurface of $M$, $\gamma$ must cross it at some time that, up to translation, we can assume to be 0. Let then $T\in\mathbb{R}$ be so that $\gamma_{T}=q$ and notice that the curve $t\mapsto\tilde{\gamma}_{t}:=\psi(\gamma_{t})\in\tilde{M}$ is a geodesic in $\tilde{M}$, thus its value $\psi(q)$ at $t=T$ is fully determined by $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{0}^{\prime}$. Since we have $\tilde{\gamma}_{0}=\tilde{\varphi}(\varphi^{-1}(\gamma_{0}))$ and, by (3.3), $\tilde{\gamma}^{\prime}_{0}=\tilde{\Phi}_{p}(\Phi_{p}^{-1}(\gamma_{0}^{\prime}))$, where $p:=\varphi^{-1}(\gamma_{0})$, we see that $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{0}^{\prime}$ only depend on $\gamma$, $\varphi$ and $\tilde{\varphi}$. In particular, they do not depend on $\psi$, proving the desired uniqueness.
We shall write $(M_{1},g_{1},\varphi_{1})\preceq(M_{2},g_{2},\varphi_{2})$, or simply $M_{1}\preceq M_{2}$, if $(M_{2},g_{2},\varphi_{2})$ extends $(M_{1},g_{1},\varphi_{1})$. The uniqueness just proved shows that this is a ‘partial order up to unique isomorphism’, i.e. that $M_{1}\preceq M_{2}$ and $M_{2}\preceq M_{1}$ imply that the maps $\psi:M_{1}\to M_{2}$ and $\tilde{\psi}:M_{2}\to M_{1}$ associated to these extensions are global causal isometric diffeomorphisms, one the inverse of the other (as by uniqueness $\tilde{\psi}\circ\psi$ must be the identity on $M_{1}$ and $\psi\circ\tilde{\psi}$ that on $M_{2}$). When this happens we write $M_{1}\cong M_{2}$.
We also notice that:
$M_{1}\preceq M_{2}\text{ and }M_{1}\not\cong M_{2}\qquad\Rightarrow\qquad M_{2}\setminus\overline{\psi(M_{1})}\text{ is not empty.}$ (3.4)
Indeed, by assumption there is $p\in M_{2}\setminus\psi(M_{1})$ and by symmetry we can assume that $p$ is in the future of $\varphi_{2}(\Sigma)$. We claim that the chronological future $I^{+}(p)$ of $p$ in $M_{2}$, that is open and not empty, does not meet $\psi(M_{1})$. Say otherwise, let $q\in I^{+}(p)\cap\psi(M_{1})$, let $\gamma:[0,1]\to M_{2}$ be a timelike curve from $p$ to $q$ and $(a,b)\subset[0,1]$ be the maximal open subinterval such that $q\in\gamma((a,b))\subset\psi(M_{1})$. Then $\gamma(a)\notin\psi(M_{1})$ and thus the curve $\psi^{-1}\circ\gamma:(a,b)\to M_{1}$ lives in the future of $\varphi_{1}(\Sigma)$, is past-inextendible and does not meet $\varphi_{1}(\Sigma)$, contradicting the assumption that $\varphi_{1}(\Sigma)$ is a Cauchy hyersurface of $M_{1}$.
The local existence and uniqueness of developments can be stated as:
Theorem 3.3.
Let $(\Sigma,h,\kappa)$ be an initial data set. Then it admits a development. Also, if $(M_{1},g_{1},\varphi_{1})$ and $(M_{2},g_{2},\varphi_{2})$ are two such developments, then they are both extension of a common development, i.e. there is a development $(M,g,\varphi)$ with $M\preceq M_{1}$ and $M\preceq M_{2}$.
In [2] the authors started from this result and, via an argument based on Zorn’s lemma, deduced existence and uniqueness of a maximal global hyperbolic development (see Theorem 3.6 below for the precise meaning of this). We are going to show that in fact for their very same proof to work Lemma 2.1 (and thus ${\rm DC}_{\omega_{1}}$) suffices.
The upper bound on the (ordinal) number of choices needed comes from the following proposition, that is a restatement of a result obtained by Geroch in [3]. We report (a minor modification of) the proof to emphasize that the signature of the non-degenerate metric tensor plays no role.
Before coming to the statement, let us agree that a smooth manifold is a Hausdorff topological space that is locally homeomorphic to $\mathbb{R}^{d}$ for some fixed $d\in\mathbb{N}$ and so that the ‘change of coordinates’ are $C^{\infty}$. Notice that we are not insisting on manifolds to be separable, and in fact
there are examples of connected smooth manifolds that are non-separable,
the typical one being the ‘long line’, see Remark 3.7 for more on this.
With this said, we have:
Proposition 3.4.
Let $M$ be a connected smooth manifold and $g$ a non-degenerate and smooth metric tensor on it.
Then $M$ is separable.
Proof.
The key of the proof is in existence of the exponential map and in its local invertibility, properties that have nothing to do with the potential lack of separability.
Since $g$ is smooth and non-degenerate, in any smooth coordinate system the Christoffel symbols $\Gamma^{k}_{ij}=\tfrac{1}{2}g^{km}(\tfrac{\partial g_{mj}}{\partial x^{i}}+\tfrac{\partial g_{im}}{\partial x^{j}}-\tfrac{\partial g_{ij}}{\partial x^{m}})$ are well defined and smooth. Hence the geodesic equation $\tfrac{{\rm d}^{2}\gamma^{k}}{{\rm d}t^{2}}+\Gamma^{k}_{ij}\frac{{\rm d}\gamma^{i}}{{\rm d}t}\frac{{\rm d}\gamma^{j}}{{\rm d}t}=0$ admits, for any initial datum, unique smooth solutions existing for some positive time. For any $x\in M$ we denote by $D_{x}\subset T_{x}M$ the domain of the exponential map, i.e. the collection of those $v\in T_{x}M$ such that there is $\gamma:[0,1]\to M$ with $\gamma_{0}=x$, $\gamma_{0}^{\prime}=v$ locally solving the geodesic equation in coordinates. Being such $\gamma$ unique, the definition $\exp_{x}(v):=\gamma_{1}$ is well posed and an argument based on smooth dependence of solutions w.r.t. initial data shows that $DM:=\{(x,v)\in TM:v\in D_{x}\subset T_{x}M\}\subset TM$ is open and that $\exp:DM\to M$ is smooth. For the same reason, the set $U_{x}\subset D_{x}$ of those $v$’s such that the differential of $\exp_{x}$ at $v$ is not singular is also open. Then the inverse function theorem grants that $\exp_{x}:U_{x}\to M$ is a local diffeomorphism and in particular that $\mathcal{U}_{x}:=\exp_{x}(U_{x})\subset M$ is open. Moreover, since $({\rm d}\exp_{x})(0)$ is, by the very definition of $\exp_{x}$, the identity, the inverse function theorem again grants that
[…]
Now fix $\bar{x}\in M$ and recursively define $M_{n}\subset M$ as: $M_{0}:=\mathcal{U}_{\bar{x}}$ and given $M_{n}$ put $M_{n+1}:=\cup_{x\in M_{n}}\mathcal{U}_{x}$. Put $M_{\infty}:=\cup_{n}M_{n}$. Since the $\mathcal{U}_{x}$’s are open, so is $M_{\infty}$. Also, since $\mathcal{U}_{\bar{x}}\subset T_{\bar{x}}M$ is separable and $\exp_{x}:U_{x}\to\mathcal{U}_{x}$ is continuous, we see that $M_{0}=\mathcal{U}_{\bar{x}}$ is separable. Then an induction argument based on the separability of $TM_{n}$ and the continuity of $\exp$ shows that $M_{n}$ is separable for every $n\in\mathbb{N}$. Hence $M_{\infty}$ is also separable.
Since $M$ is connected and $M_{\infty}\subset M$ open and separable, to conclude it suffices to prove that $M_{\infty}$ is also closed. Thus let $x\in\overline{M}_{\infty}$ and use (3.5) to find $\mathcal{V}_{x}\subset M$ with the stated properties. In particular, $\mathcal{V}_{x}\cap M_{\infty}\neq\emptyset$ and we can find $y\in V_{x}\cap M_{\infty}$ and thus $n\in\mathbb{N}$ so that $y\in\mathcal{V}_{x}\cap M_{n}$. Property (3.5) then grants that $x\in\mathcal{U}_{y}$, hence $x\in M_{n+1}\subset M_{\infty}$, as desired. ∎
Remark 3.5.
In the above statement we used regularity of $g$ to be able to define the exponential map.
We suspect the same conclusions hold even for $g$ that is merely continuous. $\blacksquare$
We turn to the main proof of the section:
Theorem 3.6.
Let $(\Sigma,h,\kappa)$ be an initial data set. Then there exists a unique maximal development $(M,g,\varphi)$ of it. This means that:
  • 1)
    (Maximality) If $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development, then $M^{\prime}\preceq M$;
  • 2)
[…]
Proof via ‘small steps’ The uniqueness claim is a direct consequence of the definition of maximality and the fact that $\preceq$ is a partial order on (isomrphism classes of) developments: if $M,\tilde{M}$ are developments satisfying $(1)$, then $\tilde{M}\preceq M$ and $M\preceq\tilde{M}$, proving that $\tilde{M}\cong M$.
The existence proof will be split in two parts. In the first part we show existence of a development $(M,g,\varphi)$ that is maximal in the sense, weaker than $(1)$, that:
  • 1’)
    (Maximality) If $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development and $M\preceq M^{\prime}$ then $M^{\prime}\cong M$.
In the second part we show that any such $M$ is also maximal in the sense of $(1)$.
We start with the first part. Let $A$ be the collection of all developments of $(\Sigma,h,\kappa)$. We avoid the usual set-theoretic issues related to the “all” above by asking the underlying set on which the manifold structures are given to be a fixed set of cardinality of continuum (notice that Proposition 3.4 ensures that this is the cardinality of any connected manifold equipped with a smooth metric tensor). Let $F\subset A$ be the collection of all the developments satisfying $(1^{\prime})$ above. Let also $\mathcal{S}$ be the collection of $\alpha$-sequences $(M_{\beta})_{\beta<\alpha}$ in $A$, $\alpha<\omega_{1}$, that are strictly increasing, meaning that for any $\beta_{1}<\beta_{2}<\alpha$ we have $M_{\beta_{1}}\preceq M_{\beta_{2}}$ and $M_{\beta_{2}}\not\preceq M_{\beta_{1}}$. We show that $A,F,\mathcal{S}$ satisfy the assumptions in Lemma 2.1.
Theorem 3.3 ensures that $\mathcal{S}$ is not empty, as it contains a 1-sequence consisting of one development. Property $(ii)$ holds trivially. $(i)$ is obvious by definition if $\alpha$ is a successor ordinal. If instead is a limit ordinal, then $M_{\alpha}$ can be built via a gluing procedure as done in [2], that requires no choice. More precisely, in [2] the following has been shown: if we are given a totally ordered collection of developments, then there exists a development bigger than, or isomorphic to, all of these. Such development is built by gluing the given ones along the maps $\psi$ coming with the relation $\preceq$. The existence of such development immediately grants that property $(i)$ holds also for $\alpha$ limit, Moreover, together with Proposition 3.4, it also implies the upper bound $(iii)$. Indeed, given a strictly increasing $\alpha$-sequence $(M_{\beta})_{\beta<\alpha}$ for some ordinal $\alpha$, possibly $\geq\omega_{1}$, what just said allows to realize, via the maps $\psi$, the $M_{\beta}$’s as subsets of a bigger development $M$ and the sets $U_{\beta}:=M_{\beta+1}\setminus\{\text{closure of $M_{\beta}$ in $M$}\}$ are open, disjoint and not empty (recall (3.4)). Proposition 3.4 ensures that $M$ is separable, thus we can only have a countable collection of such $U_{\beta}$’s, i.e. $\alpha$ is countable (alternatively, without using (3.4) we can argue as in the footnote in the introduction to conclude that the increasing $\alpha$-sequence $(M_{\beta})_{\beta<\alpha}$ of open subsets of $M$ must have countable cofinality, proving that $\alpha$ is countable).
It follows by Lemma 2.1 that $F$ is not empty, i.e. that there is a development satisfying $(1^{\prime})$, as desired.
We turn to the second part of the proof. We want to show that a development as in $(1^{\prime})$ is maximal also in the sense of $(1)$ in the statement. Inspecting the proof in [2], we see that the axiom of Choice is used, via Zorn’s lemma, to prove existence of a ‘maximal common development’ in the following sense:
[…]
In our line of thought, the existence of this maximal common development follows along the very same arguments just used: we let $A$ be the collection of developments, $F$ that of those satisfying the maximality property just stated and $\mathcal{S}$ the collection of strictly increasing $\alpha$-sequences in $A$ bounded from above by both $M_{1}$ and $M_{2}$. The non-emptiness of $\mathcal{S}$ follows by Theorem 3.3 and then that of $F$ by Lemma 2.1.
Once ‘maximal common development’ $M$ in the sense above is built, the gluing argument given in [2] shows that it is $\preceq$-maximal among all the common developments of $M_{1}$ and $M_{2}$ (not just among those comparable with $M$) and then the proof can continue as in [2]. $\square$
Proof via ‘big steps’ Let us define a way to quantify developments. Let $D\subset T\Sigma$ be countable and dense, let
$D^{T}:=\{(t,v):\ t\in\mathbb{Q},\ v\in D\text{ and }\sqrt{h(v,v)}<t\}$
and let $(t_{n},v_{n})_{n\in\mathbb{N}}$ be an enumeration of $D^{T}$. Also, let us fix an order siomorphism $\eta:[0,+\infty]\to[0,1]$.
Let $(M,g,\varphi)$ be a development of $(\Sigma,h,\kappa)$, recall that $\varphi$ induces, for any $p\in\Sigma$, an isomorphism $\Phi$ of $(\mathbb{R}\times T_{p}\Sigma,dt^{2}-h)$ and $(T_{\varphi(p)}M,g)$ sending $(1,0)$ to a future vector. For each $n\in\mathbb{N}$, $\Phi(t_{n},v_{n})$ is a future timelike vector in $M$: let $\gamma_{n}=\gamma_{n}(M)$ be the maximal geodesics (intended as solution of the geodesic equation) in $M$ with $\gamma_{n}^{\prime}(0)=\Psi(t_{n},v_{n})$. Then smoothness of $M$ grants that this is a good definition. Write $(-a_{n},b_{n})\subset\mathbb{R}$ with $a_{n},b_{n}>0$ for the interval of definition of $\gamma_{n}$. Then define
$F(M)=F(M,g,\varphi):=\sum_{n\in\mathbb{N}}\tfrac{1}{2^{n}}(\eta(b_{n})+\eta(a_{n}))\ \in[0,4].$
By definition it is obvious that $F$ is monotone,, i.e. that if $(M_{1},g_{1},\varphi_{1})$ and $(M_{2},g_{2},\varphi_{2})$ are two developments with $M_{1}\preceq M_{2}$ then $F(M_{1})\leq F(M_{2})$ (because the domain of $\gamma_{n}(M_{1})$ is contained in that of $\gamma_{n}(M_{2})$ for every $n\in\mathbb{N}$). We now observe that $F$ is strictly monotone, i.e.
$M_{1}\preceq M_{2}\text{ and }F(M_{1})=F(M_{2})\qquad\Rightarrow\qquad M_{1}\cong M_{2}.$ (3.6)
To see this we shall prove that if $M_{2}$ is a strict extension of $M_{1}$ (i.e. the canonical map $\psi:M_{1}\to M_{2}$ coming from $M_{1}\preceq M_{2}$ is not surjective), then $F(M_{2})>F(M_{1})$. By what already observed and the fact that $\eta$ is an isomorphism, to see this it suffices to find $n\in\mathbb{N}$ such that the interval of definition of $\gamma_{n}(M_{2})$ strictly contains that of $\gamma_{n}(M_{1})$.
To see this, recall by (3.4) that $M_{2}\setminus\overline{\psi(M_{1})}$ is not empty. Pick $p\in M_{2}\setminus\overline{\psi(M_{1})}$ and then an inextendible timelike geodesic $\gamma$ passing through $p$. Since $\varphi_{2}(\Sigma)$ is a Cauchy hypersurface of $M_{2}$, such $\gamma$ must intersect $\varphi_{2}(\Sigma)$. Approximating the speed of $\gamma$ at the intersection point with elements of $D^{T}$, by continuity we find $n\in\mathbb{N}$ such that the image of $\gamma_{n}(M_{2})$ intersects $M_{2}\setminus\overline{\psi(M_{1})}$. Since $\psi\circ\gamma_{n}(M_{1})$ coincides with the restriction of $\gamma_{n}(M_{2})$ to the domain of definition of $\gamma_{n}(M_{1})$, this proves that such domain of definition is strictly contained in that of $\gamma_{n}(M_{2})$, as desired.
Having established the strict monotonicity (3.6) we proceed as follows. The uniqueness property $(2)$ is obvious from the fact that $\preceq$ is a partial order. For the maximality $(1)$ we first prove the weaker claim
  • 1’)
    (Maximality) If $(M^{\prime},g^{\prime},\varphi^{\prime})$ is another development and $M\preceq M^{\prime}$ then $M^{\prime}\cong M$.
Given a development $(M,g\varphi)$ we define $G(M)\in[0,4]$ as
$G(M):=\sup\big\{F(M^{\prime})\ :\ (M^{\prime},g^{\prime},\varphi^{\prime})\text{ is a development with }M\preceq M^{\prime}\big\}$
and notice that $M_{1}\preceq M_{2}$ implies $G(M_{1})\geq G(M_{2})$. Now define recursively a sequence $(M_{n})$ of developments as follows. Let $(M_{0},g_{0},\varphi_{0})$ be an arbitrary development as given by Theorem 3.3. Having defined $(M_{n},g_{n},\varphi_{n})$ pick $(M_{n+1},g_{n+1},\varphi_{n+1})$ so that $M_{n}\preceq M_{n+1}$ and $F(M_{n+1})\geq\tfrac{1}{2}(F(M_{n})+G(M_{n}))$. The definition of $G(M)$ ensures that such $M_{n+1}$ exists. Gluing all these developments (as in [2]), we find a development $(M,g,\varphi)$ with $M_{n}\preceq M$ for every $n\in\mathbb{N}$. The construction ensures that $G(M_{n+1})-F(M_{n+1})\leq G(M_{n})-\tfrac{1}{2}(F(M_{n})+G(M_{n}))\leq G(M_{n})-F(M_{n})$ and thus $G(M_{n})-F(M_{n})\leq\tfrac{1}{2^{n}}(G(M_{0})-F(M_{0}))\to 0$, so that the monotonicities of $F,G$ imply that $G(M)=F(M)$, which by (3.6) means that $M$ is maximal as in $(1^{\prime})$.
To conclude the proof we now need to show that a development maximal as in $(1^{\prime})$ is so as in $(1)$. Arguing as in the previous proof (actually: as in [2]) this will follow if we prove the existence of a ‘maximal common development’ of two given ones. This can be achieved along the very same lines just used, by quantifying the size of common developments using $F$. We omit the details. $\square$
Remark 3.7 (The longest line).
The long line $L$ is the topological space whose underlying set is $\omega_{1}\times[0,1)$ equipped with the order topology, where the (total) order is defined by
$(\alpha,t)\leq(\beta,s)\qquad\text{whenever}\qquad\text{either $\alpha<\beta$ or ($\alpha=\beta$ and $t\leq s$)}.$
It is quite easy to see that $L\setminus\{(0,0)\}$ is locally homeomorphic to $\mathbb{R}$. A quick way to realize this is by recalling that any countable ordinal can be embedded in $\mathbb{R}$: from this fact it is immediate to see that for any point $(\bar{\alpha},\bar{t})\in L$ the set $\{(\alpha,t)\in L:(\alpha,t)\neq(0,0),\ (\alpha,t)\lneq(\bar{\alpha},\bar{t})\}$ is order isomorphic to some open interval in $\mathbb{R}$, and thus homeomorphic to such interval (as the topology on $L$ is the order topology). It is also clear that $L$ is not separable: if $((\alpha_{n},t_{n}))_{n\in\mathbb{N}}$ is any given sequence in $L$, then for $\bar{\alpha}:=\sup_{n}\alpha_{n}+1$ the open ray of points $\gneq(\bar{\alpha},0)$ does not contain any of these.
Notice that $L$ is the longest a line could be, i.e. that the following holds:
[…]
Notice that this implies that an $\rm X$} as in the statement is homeomorphic to either $\mathbb{R}$ or to $L\setminus\{0\}$ or to two copies of $L$ glued in 0.
To prove the above one line of thought is to notice that
[…]
The claim is trivial, as $U$ must be homeomorphic to an open interval, thus $\bar{U}$ is homeomorphic to either an open, or half-open or closed interval, and the conclusion is true in either case.
Given (3.8), the above claim (3.7) can be proved along the following lines. Let $U_{0}$ be an arbitrary non-empty, connected and separable open subset of $\rm X$}. As mentioned, this is homeomorphic to an interval and thus has the Lindelöf property.
We are going to recursively define $(U_{\alpha})_{\alpha<\omega_{1}}$ as in (3.7) by also ensuring that each $U_{\alpha}$ has the Lindelöf property. If $\alpha$ is a limit ordinal we put $U_{\alpha}:=\cup_{\beta<\alpha}U_{\beta}$ and notice that being a countable union of separable sets with the Lindelöf property, $U_{\alpha}$ is separable and with the Lindelöf property. On the other hand, given $U_{\alpha}$ we define $U_{\alpha+1}$ as follows. Each $p\in\overline{U}_{\alpha}$ has a neighbourhood $V_{p}$ separable and with the Lindelöf property (being homeomorphic to an interval). Also, by the Lindelöf property ensured by (3.8) there are $(p_{n})_{n\in\mathbb{N}}\subset\overline{U}_{\alpha}$ such that $\overline{U}_{\alpha}\subset\cup_{n}V_{p_{n}}$. We define $U_{\alpha+1}:=\cup_{n}V_{p_{n}}$ and notice that it is connected, separable and with the Lindelöf property.
To conclude we need to prove that $\cup_{\alpha<\omega_{1}}U_{\alpha}={\rm X}$ and since $\rm X$} is connected and such union is open , it suffices to prove that it is also closed. Let thus $p\in\overline{\cup_{\alpha<\omega_{1}}U_{\alpha}}$ and notice that since $\rm X$} is locally homeomorphic to $\mathbb{R}$ there is a sequence $(p_{n})\subset\cup_{\alpha<\omega_{1}}U_{\alpha}$ converging to $p$. Let $\alpha_{n}<\omega_{1}$ be so that $p_{n}\in U_{\alpha_{n}}$ and let $\alpha:=\sup_{n}\alpha_{n}<\omega_{1}$. We then have $p\in\overline{U}_{\alpha}\subset U_{\alpha+1}$, as desired.
It is natural to wonder whether the same result as in (3.7) holds in higher dimensions. We expect so. The argument we gave carries over to such general case with the exception of the seemingly innocent claim (3.8), that we are unable to prove. $\blacksquare$